Assume that on July 1, 2021, Togo's Sandwiches issues a $2.02 million, one-year note. Interest is payable at maturity. Determine the amount of interest expense that should be recorded in a year-end adjusting entry under each of the following independent assumptions: (Enter your answers in dollars, not in millions. Do not round intermediate calculations. Round your answers to the nearest dollar amount.)

Interest Rate Fiscal Year-End Interest Expense
1. 8 % December 31
2. 9 % September 30
3. 6 % October 31
4. 7 % January 31
Answer #1

Determine the amount of interest expense that should be recorded in a year-end adjusting entry under each of the following independent assumptions: (Enter your answers in dollars, not in millions. Do not round intermediate calculations. Round your answers to the nearest dollar amount.)

Interest Rate Fiscal Year-End Interest Expense
1. 8 % December 31 2020000*8%*6/12 = 80800
2. 9 % September 30 2020000*9%*3/12 = 45450
3. 6 % October 31 2020000*6%*4/12 = 40400
4 7 % January 31 2020000*7%*7/12 = 82483